OpenStudy (anonymous):

You are cutting vegetables for a party. You have 1/4 lb of carrots, 1/4 lb of brocolli, and 3/4 lb of peppers. How many pounds of vegetables do you have in all? A) 1 1/4 lb B) 5 lb C) 2 1/2 lb D) 1 3/4 lb Help guys!

OpenStudy (math&ing001):

1/4+1/4+3/4=5/4=1 1/4lb

OpenStudy (anonymous):

Oh thanks! I got the same answe! Would you mind to help me with another problem? I